OpenStudy (anonymous):

4+2[8+(3-1) exponent 2]

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

What are the directions please?

OpenStudy (anonymous):

solve the problem using the order of operations

OpenStudy (anonymous):

Ok start with the inner-most parentheses...(3-1)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Now will you add the 8 or use the exponent 2?

OpenStudy (anonymous):

exponent 2would be 4?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

is the answer 60

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

Now you can add the 8 with the 4

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

which is 12

OpenStudy (anonymous):

good

OpenStudy (anonymous):

4+2[8+(3-1)^2] =4+2[8+2^2] =4+2[8+4] =4+2[12] =4+24 =28

OpenStudy (anonymous):

now distribute the 2 to the 12. What do you get?

OpenStudy (anonymous):

24 + 4 which is 28 thank you all